OpenStudy (anonymous):

if the the first step in the solution of the equation -8-7x-=5-10 is "add 10",then what should the next step be?

OpenStudy (anonymous):

-8 - 7x = 5 - 10 (add 10 to both sides) -8 - 7x + 10 = 5 - 10 + 10 (combine like terms) 2 - 7x = 5 Your next step is to subtract 2 from both sides :)